Pamela is 10 years younger than Jiri. The sum of their ages is 76 . What is Jiri's age?
Answer by harpazo(655)

(x-10) + x = 76
2x-10 = 76
2x = 86
x = 86/2
x = 43
Jill is 43.
If Pamela is ten years younger, how old is Pamela?
Pamela = 43-10
Pamela = 33